OpenStudy (anonymous):

Pamela's age is 3 times Jiri's age. The sum of thier ages is 48. How old is Jiri?

OpenStudy (anonymous):

48/3

OpenStudy (anonymous):

16

OpenStudy (anonymous):

I just checked it and thats not the answer. It is quite confusing. Thank you though (:

OpenStudy (anonymous):

P + J = 64 (given in problem) J = 64- P (solving for J) P = 3J (given in problem) P = 3(64-P) (substituting step one solution into equation) P = 192-3P (distributive law of multiplication) 4P = 192 (adding 3P to both sides) P = 48 (divide both sides by 4) J = 64 - 48 (substituting our solution of P in with step 2, to solve for J J = 16 (simplifying the expression) CHECK Does J = three times P 16 x 3 = 48. Yes. Does J + P = 64? 16+48 = 64. Yes!

OpenStudy (anonymous):

example

OpenStudy (triciaal):